USE EXCEL TO CALCULATE THE FREQUENCIES AS SHOWN BELOW. PLEASE PROVIDE EXCEL FORMULA USED.

Frequency Distribution
Low High Bins Frequency
-67.0 -56.6 (-67, -56.6]
-56.6 -46.2 (-56.6, -46.2]
-46.2 -35.8 (-46.2, -35.8]
-35.8 -25.4 (-35.8, -25.4]
-25.4 -15.0 (-25.4, -15]
-15.0 -4.6 (-15, -4.6]
-4.6 5.8 (-4.6, 5.8]
5.8 16.2 (5.8, 16.2]
16.2 26.6 (16.2, 26.6]
26.6 37.0 (26.6, 37]
37.0 47.4 (37, 47.4]
47.4 57.8 (47.4, 57.8]
57.8 68.2 (57.8, 68.2]
68.2 78.6 (68.2, 78.6]
78.6 89.0 (78.6, 89]
89.0 99.4 (89, 99.4]
2.0
-61.0
-9.0
14.0
43.0
-16.0
-11.0
8.0
20.0
-15.0
0.0
-37.0
9.0
14.0
40.0
-29.0
38.0
-49.0
65.0
27.0
89.0
-16.0
83.0
72.0
-3.0
-21.0
51.0
36.0
26.0
-45.0
-6.0
14.0
-3.0
55.0
39.0
-6.0
-35.0
25.0
27.0
6.0
12.0
37.0
-12.0
35.0
24.0
15.0
35.0
-67.0
2.0
-32.0
-34.0
-32.0
22.0
51.0
33.0
10.0
-28.0
2.0
12.0
17.0
44.0
-29.0
22.0
-28.0
-18.0
-12.0
-19.0
11.0
62.0
50.0
-27.0
46.0
62.0
-26.0
-54.0
-61.0
45.0
-51.0
-41.0
14.0
9.0
-56.0
0.0
-34.0
-52.0
-28.0
14.0
63.0
23.0
-46.0
29.0
1.0
1.0
20.0
1.0
25.0
-4.0
29.0
-15.0
78.0
Answer #1

The answer is:

Frequency Distribution
Low High Bins Frequency
-67.0 -56.6 (-67, -56.6] 3
-56.6 -46.2 (-56.6, -46.2] 5
-46.2 -35.8 (-46.2, -35.8] 4
-35.8 -25.4 (-35.8, -25.4] 12
-25.4 -15.0 (-25.4, -15] 5
-15.0 -4.6 (-15, -4.6] 8
-4.6 5.8 (-4.6, 5.8] 11
5.8 16.2 (5.8, 16.2] 14
16.2 26.6 (16.2, 26.6] 10
26.6 37.0 (26.6, 37] 8
37.0 47.4 (37, 47.4] 8
47.4 57.8 (47.4, 57.8] 4
57.8 68.2 (57.8, 68.2] 4
68.2 78.6 (68.2, 78.6] 2
78.6 89.0 (78.6, 89] 1
89.0 99.4 (89, 99.4] 1
Total 100

The formulas are:

Frequency Distribution
Low High Bins Frequency
-67.0 -56.6 (-67, -56.6] =COUNTIFS(A1:A100,">-68",A1:A100,"<-56.6")
-56.6 -46.2 (-56.6, -46.2] =COUNTIFS(A1:A100,">-56.7",A1:A100,"<-46.2")
-46.2 -35.8 (-46.2, -35.8] =COUNTIFS(A1:A100,">-46.3",A1:A100,"<-35.8")
-35.8 -25.4 (-35.8, -25.4] =COUNTIFS(A1:A100,">-35.9",A1:A100,"<-25.4")
-25.4 -15.0 (-25.4, -15] =COUNTIFS(A1:A100,">-25.5",A1:A100,"<-15")
-15.0 -4.6 (-15, -4.6] =COUNTIFS(A1:A100,">-15.1",A1:A100,"<-4.6")
-4.6 5.8 (-4.6, 5.8] =COUNTIFS(A1:A100,">-4.7",A1:A100,"<5.8")
5.8 16.2 (5.8, 16.2] =COUNTIFS(A1:A100,">5.7",A1:A100,"<16.2")
16.2 26.6 (16.2, 26.6] =COUNTIFS(A1:A100,">16.1",A1:A100,"<26.6")
26.6 37.0 (26.6, 37] =COUNTIFS(A1:A100,">26.5",A1:A100,"<37")
37.0 47.4 (37, 47.4] =COUNTIFS(A1:A100,">36.9",A1:A100,"<47.4")
47.4 57.8 (47.4, 57.8] =COUNTIFS(A1:A100,">47.3",A1:A100,"<57.8")
57.8 68.2 (57.8, 68.2] =COUNTIFS(A1:A100,">57.7",A1:A100,"<68.2")
68.2 78.6 (68.2, 78.6] =COUNTIFS(A1:A100,">68.1",A1:A100,"<78.6")
78.6 89.0 (78.6, 89] =COUNTIFS(A1:A100,">78.5",A1:A100,"<89")
89.0 99.4 (89, 99.4] =COUNTIFS(A1:A100,">88.9",A1:A100,"<99.4")
Total =SUM(G3:G18)

The Excel output is:

The Excel formulas are:

Please give me a thumbs-up if this helps you out. Thank you!